# DESI results and Dark Energy from QCD topological sectors

### v1

@@ Abstract
We present a physically motivated model of dark energy (DE) rooted in the topological structure of the Quantum ChromoDynamic (QCD) vacuum. In this framework, DE emerges as the Universe expands, from the difference in vacuum energy between the expanding Universe and Minkowski spacetime, driven by QCD vacuum topological sectors. This leads to a modification of the dark energy term in the Friedmann equation, which then scales with the Hubble parameter as $\rho_{\rm DE}(t)\propto H(t)$ when dark energy dominates the expansion of the Universe. The QCD scale, $\Lambda_{\rm QCD}\sim 100~{}{\rm MeV}$, naturally sets the energy density of DE and provides a compelling explanation for why its impact on cosmic expansion becomes significant only in the recent cosmological epoch. Importantly, the entire framework is grounded in Standard Model (SM) physics, involving no new fields or coupling constants. Key predictions of the model include: (a) A present-day DE equation of state $w_{\rm DE}>-1$, asymptotically approaching the de Sitter limit $w_{\rm DE}=-1$ in the future, with the corresponding asymptotic Hubble constant $\overline{H}$ set by $\Lambda_{\rm QCD}$. (b) At redshifts $z\geq 0$, $w_{\rm DE}$ can lie above or below $-1$ and may cross this boundary multiple times—behavior qualitatively consistent with recent DESI observations. (c) The solution to the Friedmann equation in this framework can deviate from the canonical $\Lambda$CDM form at $z\geq 0$. (d) If such deviation occurs, it can be tested using cosmological observations, CMB anisotropies, BAO, SNIa, and large-scale structure, which we propose to explore in future work. (e) Finally, we point out a potential connection between our framework and the observed $H_{0}$ tension.
@@ I Introduction
Since the discovery of the accelerated expansion rate of the Universe interpreted as a mysterious dark energy (DE) [1, 2] a countless number of models have been proposed to explain its existence. The recent DESI results are the most precise measurement of the DE equation of state so far [3]. If these results are confirmed, explaining simultaneously $w_{\rm DE}>-1$ today and a redshift evolution such that $w_{\rm DE}(z>0)<-1$ will represent a real challenge for DE models.
The possibility that $w_{\rm DE}$ may vary with redshift came as a surprise, although there are theoretical arguments suggesting that such a dependence is to be expected [4]. Even more unexpected is the possibility that $w_{\rm DE}$ could cross the $w_{\rm DE}=-1$ line. The regime where $w_{\rm DE}<-1$—commonly referred to as the phantom regime—is already highly puzzling, as it leads to violations of unitarity and causality when formulated within the framework of quantum field theory (QFT). The additional feature of crossing the $w_{\rm DE}=-1$ boundary—known as quintom behavior—further challenges fundamental principles of QFT, as discussed in the review by [5]. For a recent summary of this topic, including numerous original references and context in light of the DESI results, we refer the reader to the short overview in [6]. The origin of all the issues associated with the $w_{\rm DE}\leq-1$ behavior lies in the presence of a dynamical canonical field $\phi$ governed by a potential $V(\phi)$. In such cases, the theory cannot be formulated consistently due to internal instabilities and related problems, including a negative square of the speed of sound, $c_{s}^{2}<0$, among other fundamental inconsistencies.
In this paper, we develop a dramatically different approach based on decade old idea, proposed by [7, 8, 9], which finds its root in a 1967 paper by Zeldovich [10]. Essentially, the Zeldovich’s prescription (expressed in modern terms) can be formulated as follows. One must compute the non-perturbative vacuum energy in Quantum Chromodynamics (QCD) in an expanding universe, characterized by dimensional parameters such as the QCD scale $\Lambda_{\mathrm{QCD}}$ and the Hubble constant $H$. A similar computation should then be performed for the QCD vacuum energy in Minkowski spacetime. The difference between these two results—referred to as the leftover in this work—enters the right-hand side of the Einstein equation, in accordance with Zeldovich’s prescription.
@@ II The topology as the source of the gravitating vacuum energy
The objective of this section is to provide an overview of the basic ideas proposed in [7, 8, 9] concerning the nature of vacuum energy in strongly coupled gauge theories such as quantum chromodynamics (QCD). In this framework, the Universe will naturally evolve towards the de Sitter geometry, as a result of the presence of topologically nontrivial sectors $|k\rangle$ in a time-dependent, expanding universe In our proposal, the Dark Energy (DE) is identified with the vacuum energy produced by tunnelling transitions between different topological sectors. The amplitude of DE is associated with the QCD energy scale, while its time variation arises from changes in the topological sectors $|k\rangle$ and the corresponding tunnelling transition rates in an expanding universe.
@@ II.1 Intuitive picture
Consider, for example, the Aharonov-Casher effect [11]: an external charge is inserted into a superconductor where the electric field is exponentially screened, $\sim\exp(-r/\lambda)$, with $\lambda$ being the penetration depth. Despite the very short penetration length, a neutral magnetic fluxon remains sensitive to the external charge even at arbitrarily large distances $r\gg\lambda$, despite the screening of the physical field (which is equivalent to the presence of a mass gap in our system). This genuine quantum effect is purely topological and non-local in nature, and it can be understood in terms of the dynamics of the gauge sectors responsible for the long-range behavior.
@@
We conclude this discussion on the nature of Dark Energy (DE)—and its sensitivity to arbitrarily large distances despite the presence of a mass gap in QCD—by noting that the existence of long-range forces in gauge theories with a gap has been suspected for a long time. This idea appears, for instance, in a paper by Lüscher [12] titled ”The secret long range force in quantum field theories with instantons.” The technical origin of these long-range forces can be traced back to an earlier work by ’t Hooft [13], where the computation of the instanton density requires introducing an infrared (IR) cutoff to properly account for zero gauge modes. A crucial aspect of this calculation is that the correction due to the IR cutoff is only power suppressed—not exponentially suppressed, as one might naively expect from the presence of a gap in the system. See also [14] for several relevant comments on this point. The same long-range interaction can be described in terms of the so-called Veneziano ghost [15]; see Appendix A. In exactly solvable models, one can explicitly see that the Veneziano ghost is in fact an auxiliary, topological, non-propagating field—commonly used in condensed matter physics to describe topologically ordered systems—which lacks a canonical kinetic term. We briefly review the connections between these various descriptions of this highly nontrivial phenomenon with specific references in Appendix A.
@@ II.2 Generating DE from QCD topological sectors
In the approach of [7, 8, 9], the vacuum energy entering the Friedmann equation is defined as $\Delta\rho\equiv\rho_{\rm FRW}-\rho_{\mathrm{Mink}}$, where $\rho_{\mathrm{Mink}}$ represents the vacuum energy in Minkowski spacetime. This definition of vacuum energy was first proposed in 1967 by Zeldovich [10], who argued that $\rho_{\text{vac}}=\Delta\rho\sim Gm_{p}^{6}$ must be proportional to the gravitational constant, with $m_{p}$ being the proton mass (i.e., the QCD scale). In the following decades, various papers, using the same definition for $\Delta\rho$ in Einstein’s field equations, have been written by researchers in different fields, including particle physics, cosmology, condensed matter physics, see [8] for the references and details.
The computation of $\Delta\rho$ for arbitrary geometries is currently infeasible due to several technical challenges (see [8] for a detailed discussion). However, in certain special cases, such calculations have been successfully performed and form a key part of the theoretical framework developed in [7, 8, 9]. One such case involves the relativistic hyperbolic spacetime $\mathbb{H}^{3}_{\kappa}\times\mathbb{S}^{1}_{\kappa^{-1}}$, characterized by the curvature parameter $\kappa$, where $\mathbb{H}^{3}_{\kappa}$ denotes three-dimensional hyperbolic space, and $\mathbb{S}^{1}_{\kappa^{-1}}$ is a one-dimensional circle (or ring) with circumference $\kappa^{-1}$. The vacuum energy associated with this geometry is denoted by $E_{\rm vac}[\mathbb{H}^{3}_{\kappa}\times\mathbb{S}^{1}_{\kappa^{-1}}]$. In comparison, the vacuum energy of flat Minkowski spacetime is given by $E_{\rm vac}[\mathbb{R}^{3}\times\mathbb{S}^{1}]\sim\Lambda_{\rm QCD}^{4}$, where $\mathbb{R}^{3}$ is three-dimensional Euclidean space and $\mathbb{S}^{1}$ is a one-dimensional circle. According to the Zeldovich prescription as described above, the central claim of [8] is that the vacuum energy difference between these two geometries, $\Delta E_{\rm vac}$, receives a linear correction proportional to $\kappa$, i.e.:
@@
where numerical factors are omitted and $c_{\kappa}$ is a dimensionless numerical coefficient of order one. The crucial minus sign ()(-)( - ) is retained, which shows that the vacuum energy difference is positive. The negative sign associated with the QCD vacuum energy is a well-known feature of QCD, originally introduced through the so-called Bag constant, a phenomenological parameter in the MIT bag model [16, 17]. This Bag constant was later shown to be expressible in terms of the energy-momentum tensor and its trace anomaly. The negative sign, proportional to $c_{\kappa}\kappa$, arises from explicit computations [8], and can be interpreted as the negative sign typically encountered in Casimir-type energy calculations for systems with boundary conditions or constraints, in contrast to Minkowski space-time.
For de Sitter spacetime, it is not currently possible to perform analogous computations [8]. One of the main obstacles is that monopole solutions with nontrivial holonomy contributing to $\Delta E_{\rm vac}$ —similar to those defined on $\mathbb{H}^{3}_{\kappa}\times\mathbb{S}^{1}_{\kappa^{-1}}$— are not known in the context of de Sitter spacetime. However, as argued in [9], one can conjecture that the resulting expression is expected to closely resemble Eq.(1). It means that, in de Sitter spacetime, holonomy is expected to emerge dynamically (as the Universe expands), and the role of $\kappa$ in Eq.(1) is assumed by the Hubble parameter in the de Sitter Universe, with the replacement $c_{\kappa}\sim 1$ by $c_{H}\sim 1$.
@@
3. The relevant topological Euclidean configurations that saturate the vacuum energy (1) can be interpreted as three-dimensional magnetic monopoles wrapping around the $\mathbb{S}^{1}$ direction [8]. These configurations are characterized by non-vanishing holonomy —a measure of the gauge field’s behavior around a closed loop—, which ultimately leads to a linear (rather than quadratic) correction $\sim\kappa$ to the vacuum energy density. Additional discussion of this linear dependence can be found in Appendix B with references on the original results. This entire gauge configuration represents only a saddle point in the Euclidean path integral—a mathematical construction used to describe quantum tunnelling events. It does not correspond to the propagation of a physical degree of freedom capable of transmitting information or signals. In a cosmological context, such configurations are highly unconventional: they inherently describe non-local physics, since the holonomy is itself a non-local quantity. This non-locality is precisely why these effects cannot be captured by any local quantum field theory, as noted in item 2. It is a similar situation to the Aharonov-Casher effect, which was discussed in Section II.1 to provide an intuitive picture.
4. Equation (1) is consistent with earlier findings in the weakly coupled “deformed QCD” model, where all computations are analytical. In that model, the sensitivity of the vacuum energy to very large distances can be studied by enclosing the system in a box of size $\mathbb{L}$. As shown in [14], the corrections to the vacuum energy scale linearly with the inverse size, $\sim\mathbb{L}^{-1}$—a behavior analogous to the role played by the parameter $\kappa$ in equation (1). This model closely resembles the system considered in the present work, since the vacuum energy in the deformed QCD model is also dominated by monopoles with nontrivial holonomy and features a mass gap. In contrast, conventional ’t Hooft instantons with trivial holonomy produce only quadratic corrections, $\sim\mathbb{L}^{-2}$, as also noted in [14].
5. Equation (1) is also consistent with lattice simulations presented in [18], where the author investigates the rate of particle production in a de Sitter background. The results show that the production rate is linearly proportional to the Hubble constant, scaling as $\sim H$, rather than the expected $H^{2}$.
@@ II.3 The Equation of State (EoS) for QCD-induced DE in de Sitter case
This corresponds precisely to the estimate originally proposed by Zeldovich long ago [10], provided one replaces $m_{p}\rightarrow{\Lambda_{\mathrm{QCD}}}$ in his formula. Taking $\Lambda_{\mathrm{QCD}}\approx 100\;{\rm MeV}$ and assuming that the dimensionless numerical coefficient $c_{H}$ remains constant [^4] We do not lose any generality by fixing $\Lambda_{\mathrm{QCD}}\approx 100\;{\rm MeV}$. This is because the dimensionless parameter $c_{H}$ can always be redefined to absorb all numerical coefficients that arise in the calculations. In particular, a small numerical QCD-related factor $\propto m_{q}/\Lambda_{\mathrm{QCD}}\approx 0.05$, which consistently accompanies tunnelling transitions in QCD, is also absorbed into $c_{H}$. In this estimate, the quark mass is taken to be approximately $m_{q}\approx 5$ MeV.:
@@
which are indeed very close to the observed values today [^5] It is also worth mentioning that the numerical coincidence between the observed value of $\rho_{\rm DE}$ and the estimate in equation (3) was the primary motivation for the proposal in [19, 20], which suggested that the driving force behind dark energy is the nontrivial dynamics of the topological sectors in strongly coupled QCD—originally formulated in terms of the Veneziano ghost (albeit without a clear understanding at the time of the physical basis for the formula). It took several years before the key elements outlined in items 1–5 of Section II.2 were fully understood. The connection between the Veneziano ghost and the topological auxiliary field was also elucidated later in [21]. Additional remarks on this relationship are provided in Appendix A..
@@
The main results of this subsection can be summarized as follows: The pure de Sitter state can be characterized by a single parameter—the Hubble constant $\overline{H}$—and the tunneling transitions in QCD generate dark energy with the same characteristics as a cosmological constant, as given by (6) and its equation of state (7). We emphasize that no ad hoc constants are introduced into the system. In principle, the constant $c_{H}$ could be computed from first principles (although this is not currently feasible, as discussed in [8]), similar to model computations briefly reviewed in Sect. II.2. This constant is dynamically generated, as Zeldovich conjectured more than half a century ago [10], if one replaces $m_{p}\rightarrow\Lambda_{\mathrm{QCD}}$ in his formula. What happens if we slowly vary ${H}$ over time? According to the adiabatic theorem, the relations (6) and (7) are expected to hold as long as the de Sitter state continues to dominate the cosmic evolution. We will make use of this feature in the next Sect. III, where we investigate deviations from exact de Sitter behavior (7) with the aim of comparing to recent DESI results.

### v2

@@ Abstract
[…]
show]waerbeke@phas.ubc.ca
@@ I Introduction
Since the discovery of the accelerated expansion rate of the Universe interpreted as a mysterious dark energy (DE) Perlmutter et al. (1999); Riess et al. (1998) a countless number of models have been proposed to explain its existence. The recent DESI results are the most precise measurement of the DE equation of state so far DESI Collaboration et al. (2025). If these results are confirmed, explaining simultaneously $w_{\rm DE}>-1$ today and a redshift evolution such that $w_{\rm DE}(z>0)<-1$ will represent a real challenge for DE models.
The possibility that $w_{\rm DE}$ may vary with redshift came as a surprise, although there are theoretical arguments suggesting that such a dependence is to be expected Brandenberger (2025). Even more unexpected is the possibility that $w_{\rm DE}$ could cross the $w_{\rm DE}=-1$ line. The regime where $w_{\rm DE}<-1$—commonly referred to as the phantom regime—is already highly puzzling, as it leads to violations of unitarity and causality when formulated within the framework of quantum field theory (QFT). The additional feature of crossing the $w_{\rm DE}=-1$ boundary—known as quintom behavior—further challenges fundamental principles of QFT, as discussed in the review by Cai et al. (2010). For a recent summary of this topic, including numerous original references and context in light of the DESI results, we refer the reader to the short overview in Cai et al. (2025). The origin of all the issues associated with the $w_{\rm DE}\leq-1$ behavior lies in the presence of a dynamical canonical field $\phi$ governed by a potential $V(\phi)$. In such cases, the theory cannot be formulated consistently due to internal instabilities and related problems, including a negative square of the speed of sound, $c_{s}^{2}<0$, among other fundamental inconsistencies.
In this paper, we develop a dramatically different approach based on decade old idea, proposed by Zhitnitsky (2014, 2015); Barvinsky and Zhitnitsky (2018), which finds its root in a 1967 paper by Zeldovich Zeldovich (1967). Essentially, the Zeldovich’s prescription (expressed in modern terms) can be formulated as follows. One must compute the non-perturbative vacuum energy in Quantum Chromodynamics (QCD) in an expanding universe, characterized by dimensional parameters such as the QCD scale $\Lambda_{\mathrm{QCD}}$ and the Hubble constant $H$. A similar computation should then be performed for the QCD vacuum energy in Minkowski spacetime. The difference between these two results—referred to as the leftover in this work—enters the right-hand side of the Einstein equation, in accordance with Zeldovich’s prescription.
@@ II The topology as the source of the gravitating vacuum energy
The objective of this section is to provide an overview of the basic ideas proposed in Zhitnitsky (2014, 2015); Barvinsky and Zhitnitsky (2018) concerning the nature of vacuum energy in strongly coupled gauge theories such as quantum chromodynamics (QCD). In this framework, the Universe will naturally evolve towards the de Sitter geometry, as a result of the presence of topologically nontrivial sectors $|k\rangle$ in a time-dependent, expanding universe In our proposal, the Dark Energy (DE) is identified with the vacuum energy produced by tunnelling transitions between different topological sectors. The amplitude of DE is associated with the QCD energy scale, while its time variation arises from changes in the topological sectors $|k\rangle$ and the corresponding tunnelling transition rates in an expanding universe.
@@ II.1 Intuitive picture
Consider, for example, the Aharonov-Casher effect Reznik and Aharonov (1989): an external charge is inserted into a superconductor where the electric field is exponentially screened, $\sim\exp(-r/\lambda)$, with $\lambda$ being the penetration depth. Despite the very short penetration length, a neutral magnetic fluxon remains sensitive to the external charge even at arbitrarily large distances $r\gg\lambda$, despite the screening of the physical field (which is equivalent to the presence of a mass gap in our system). This genuine quantum effect is purely topological and non-local in nature, and it can be understood in terms of the dynamics of the gauge sectors responsible for the long-range behavior.
@@
We conclude this discussion on the nature of Dark Energy (DE)—and its sensitivity to arbitrarily large distances despite the presence of a mass gap in QCD—by noting that the existence of long-range forces in gauge theories with a gap has been suspected for a long time. This idea appears, for instance, in a paper by Lüscher Luscher (1978) titled ”The secret long range force in quantum field theories with instantons.” The technical origin of these long-range forces can be traced back to an earlier work by ’t Hooft ’t Hooft (1976), where the computation of the instanton density requires introducing an infrared (IR) cutoff to properly account for zero gauge modes. A crucial aspect of this calculation is that the correction due to the IR cutoff is only power suppressed—not exponentially suppressed, as one might naively expect from the presence of a gap in the system. See also Thomas and Zhitnitsky (2012) for several relevant comments on this point. The same long-range interaction can be described in terms of the so-called Veneziano ghost Veneziano (1979); see Appendix A. In exactly solvable models, one can explicitly see that the Veneziano ghost is in fact an auxiliary, topological, non-propagating field—commonly used in condensed matter physics to describe topologically ordered systems—which lacks a canonical kinetic term. We briefly review the connections between these various descriptions of this highly nontrivial phenomenon with specific references in Appendix A.
@@ II.2 Generating DE from QCD topological sectors
In the approach of Zhitnitsky (2014, 2015); Barvinsky and Zhitnitsky (2018), the vacuum energy entering the Friedmann equation is defined as $\Delta\rho\equiv\rho_{\rm FRW}-\rho_{\mathrm{Mink}}$, where $\rho_{\mathrm{Mink}}$ represents the vacuum energy in Minkowski spacetime. This definition of vacuum energy was first proposed in 1967 by Zeldovich Zeldovich (1967), who argued that $\rho_{\text{vac}}=\Delta\rho\sim Gm_{p}^{6}$ must be proportional to the gravitational constant, with $m_{p}$ being the proton mass (i.e., the QCD scale). In the following decades, various papers, using the same definition for $\Delta\rho$ in Einstein’s field equations, have been written by researchers in different fields, including particle physics, cosmology, condensed matter physics, see Zhitnitsky (2015) for the references and details.
The computation of $\Delta\rho$ for arbitrary geometries is currently infeasible due to several technical challenges (see Zhitnitsky (2015) for a detailed discussion). However, in certain special cases, such calculations have been successfully performed and form a key part of the theoretical framework developed in Zhitnitsky (2014, 2015); Barvinsky and Zhitnitsky (2018). One such case involves the relativistic hyperbolic spacetime $\mathbb{H}^{3}_{\kappa}\times\mathbb{S}^{1}_{\kappa^{-1}}$, characterized by the curvature parameter $\kappa$, where $\mathbb{H}^{3}_{\kappa}$ denotes three-dimensional hyperbolic space, and $\mathbb{S}^{1}_{\kappa^{-1}}$ is a one-dimensional circle (or ring) with circumference $\kappa^{-1}$. The vacuum energy associated with this geometry is denoted by $E_{\rm vac}[\mathbb{H}^{3}_{\kappa}\times\mathbb{S}^{1}_{\kappa^{-1}}]$. In comparison, the vacuum energy of flat Minkowski spacetime is given by $E_{\rm vac}[\mathbb{R}^{3}\times\mathbb{S}^{1}]\sim\Lambda_{\rm QCD}^{4}$, where $\mathbb{R}^{3}$ is three-dimensional Euclidean space and $\mathbb{S}^{1}$ is a one-dimensional circle. According to the Zeldovich prescription as described above, the central claim of Zhitnitsky (2015) is that the vacuum energy difference between these two geometries, $\Delta E_{\rm vac}$, receives a linear correction proportional to $\kappa$, i.e.:
@@
where numerical factors are omitted and $c_{\kappa}$ is a dimensionless numerical coefficient of order one. The crucial minus sign ()(-) is retained, which shows that the vacuum energy difference is positive. The negative sign associated with the QCD vacuum energy is a well-known feature of QCD, originally introduced through the so-called Bag constant, a phenomenological parameter in the MIT bag model Chodos et al. (1974a, b). This Bag constant was later shown to be expressible in terms of the energy-momentum tensor and its trace anomaly. The negative sign, proportional to $c_{\kappa}\kappa$, arises from explicit computations Zhitnitsky (2015), and can be interpreted as the negative sign typically encountered in Casimir-type energy calculations for systems with boundary conditions or constraints, in contrast to Minkowski space-time.
For de Sitter spacetime, it is not currently possible to perform analogous computations Zhitnitsky (2015). One of the main obstacles is that monopole solutions with nontrivial holonomy contributing to $\Delta E_{\rm vac}$ —similar to those defined on $\mathbb{H}^{3}_{\kappa}\times\mathbb{S}^{1}_{\kappa^{-1}}$— are not known in the context of de Sitter spacetime. However, as argued in Barvinsky and Zhitnitsky (2018), one can conjecture that the resulting expression is expected to closely resemble Eq.(1). It means that, in de Sitter spacetime, holonomy is expected to emerge dynamically (as the Universe expands), and the role of $\kappa$ in Eq.(1) is assumed by the Hubble parameter in the de Sitter Universe, with the replacement $c_{\kappa}\sim 1$ by $c_{H}\sim 1$.
@@
3. The relevant topological Euclidean configurations that saturate the vacuum energy (1) can be interpreted as three-dimensional magnetic monopoles wrapping around the $\mathbb{S}^{1}$ direction Zhitnitsky (2015). These configurations are characterized by non-vanishing holonomy —a measure of the gauge field’s behavior around a closed loop—, which ultimately leads to a linear (rather than quadratic) correction $\sim\kappa$ to the vacuum energy density. Additional discussion of this linear dependence can be found in Appendix B with references on the original results. This entire gauge configuration represents only a saddle point in the Euclidean path integral—a mathematical construction used to describe quantum tunnelling events. It does not correspond to the propagation of a physical degree of freedom capable of transmitting information or signals. In a cosmological context, such configurations are highly unconventional: they inherently describe non-local physics, since the holonomy is itself a non-local quantity. This non-locality is precisely why these effects cannot be captured by any local quantum field theory, as noted in item 2. It is a similar situation to the Aharonov-Casher effect, which was discussed in Section II.1 to provide an intuitive picture.
4. Equation (1) is consistent with earlier findings in the weakly coupled “deformed QCD” model, where all computations are analytical. In that model, the sensitivity of the vacuum energy to very large distances can be studied by enclosing the system in a box of size $\mathbb{L}$. As shown in Thomas and Zhitnitsky (2012), the corrections to the vacuum energy scale linearly with the inverse size, $\sim\mathbb{L}^{-1}$—a behavior analogous to the role played by the parameter $\kappa$ in equation (1). This model closely resembles the system considered in the present work, since the vacuum energy in the deformed QCD model is also dominated by monopoles with nontrivial holonomy and features a mass gap. In contrast, conventional ’t Hooft instantons with trivial holonomy produce only quadratic corrections, $\sim\mathbb{L}^{-2}$, as also noted in Thomas and Zhitnitsky (2012).
5. Equation (1) is also consistent with lattice simulations presented in Yamamoto (2014), where the author investigates the rate of particle production in a de Sitter background. The results show that the production rate is linearly proportional to the Hubble constant, scaling as $\sim H$, rather than the expected $H^{2}$.
@@ II.3 The Equation of State (EoS) for QCD-induced DE in de Sitter case
This corresponds precisely to the estimate originally proposed by Zeldovich long ago Zeldovich (1967), provided one replaces $m_{p}\rightarrow{\Lambda_{\mathrm{QCD}}}$ in his formula. Taking $\Lambda_{\mathrm{QCD}}\approx 100\;{\rm MeV}$ and assuming that the dimensionless numerical coefficient $c_{H}$ remains constant [^4] We do not lose any generality by fixing $\Lambda_{\mathrm{QCD}}\approx 100\;{\rm MeV}$. This is because the dimensionless parameter $c_{H}$ can always be redefined to absorb all numerical coefficients that arise in the calculations. In particular, a small numerical QCD-related factor $\propto m_{q}/\Lambda_{\mathrm{QCD}}\approx 0.05$, which consistently accompanies tunnelling transitions in QCD, is also absorbed into $c_{H}$. In this estimate, the quark mass is taken to be approximately $m_{q}\approx 5$ MeV.:
@@
which are indeed very close to the observed values today [^5] It is also worth mentioning that the numerical coincidence between the observed value of $\rho_{\rm DE}$ and the estimate in equation (3) was the primary motivation for the proposal in Urban and Zhitnitsky (2010a, b), which suggested that the driving force behind dark energy is the nontrivial dynamics of the topological sectors in strongly coupled QCD—originally formulated in terms of the Veneziano ghost (albeit without a clear understanding at the time of the physical basis for the formula). It took several years before the key elements outlined in items 1–5 of Section II.2 were fully understood. The connection between the Veneziano ghost and the topological auxiliary field was also elucidated later in Zhitnitsky (2013). Additional remarks on this relationship are provided in Appendix A..
@@
The main results of this subsection can be summarized as follows: The pure de Sitter state can be characterized by a single parameter—the Hubble constant $\overline{H}$—and the tunnelling transitions in QCD generate dark energy with the same characteristics as a cosmological constant, as given by (6) and its equation of state (7). We emphasize that no ad hoc constants are introduced into the system. In principle, the constant $c_{H}$ could be computed from first principles (although this is not currently feasible, as discussed in Zhitnitsky (2015)), similar to model computations briefly reviewed in Sect. II.2. This constant is dynamically generated, as Zeldovich conjectured more than half a century ago Zeldovich (1967), if one replaces $m_{p}\rightarrow\Lambda_{\mathrm{QCD}}$ in his formula. What happens if we slowly vary H{H} over time? According to the adiabatic theorem, the relations (6) and (7) are expected to hold as long as the de Sitter state continues to dominate the cosmic evolution. We will make use of this feature in the next Sect. III, where we investigate deviations from exact de Sitter behavior (7) with the aim of comparing to recent DESI results.